# Harmonics of Learning: Universal Fourier Features Emerge in Invariant Networks

### v2

@@ Appendix B Spectral Invariants
In this section, we overview the theory of invariants over $\langle G\rangle$, i.e. (polynomial) maps $\langle G\rangle\rightarrow\mathbb{C}$ that are invariant with respect to the action by $G$. To this end, we recall the following notion.
Definition 10.
@@
$\beta_{\underline{\rho}}(x)=\hat{x}_{\rho_{1}}\cdots\hat{x}_{\rho_{n}}\overline{\hat{x}}_{\rho_{1}\cdots\rho_{n}}$ (45)
The spectra of order $n$ are invariant polynomials of degree $n+1$ containing one conjugate variable. The presence of the latter is necessary for invariance. For $n=1,2$ they are alternatively referred to as power spectra and bispectra respectively. Note that the power spectra reduce simply to $\beta_{\rho}(x)=|\hat{x}_{\rho}|^{2}$, $\rho\in G^{\vee}$, and constitute a standard tool in signal processing. Bispectra, together with higher-order spectra, were first introduced in [27]. It is immediate to see that spectra generate all the polynomial invariants of $\langle G\rangle$ (see also [48, Theorem 2.1.4]).
@@
Proof.
This follows from interpreting the invariance condition via the Fourier transform. Namely, given $\underline{\rho}\in(G^{\vee})^{n+1}$ consider the monomial over $\langle G^{\vee}\rangle$ defined by $\hat{x}_{\rho_{1}}\cdots\hat{x}_{\rho_{n}}\overline{\hat{x}}_{\rho_{n+1}}$. Since $g\cdot\hat{x}=(\overline{\rho(g)}\ \hat{x}_{\lambda})_{\lambda\in G^{\vee}}$, the monomial is invariant if and only if $\rho_{1}(g)\cdots\rho_{n}(g)\overline{\rho}_{n+1}(g)=1$ for all $g\in G$, i.e. $\rho_{n+1}=\rho_{1}\cdots\rho_{n}$. Since monomials linearly generate polynomials, the claim follows. ∎
A remarkable aspect of spectra of even order is the fact that they jointly determine real (generic) elements of $\langle G\rangle$ up to the action by $G$ – a property known as completeness. This was first shown in [45]. For convenience, we report below a simple proof for finite commutative groups.
@@
Proof.
By setting $\underline{\rho}=(1,\cdots,1)$ we see that $\langle 1,x\rangle^{n+1}=\langle 1,y\rangle^{n+1}\in\mathbb{R}\setminus\{0\}$ and therefore $\langle 1,x\rangle=\langle 1,y\rangle$ since $n$ is even. For $\rho\in G^{\vee}$, by setting $\underline{\rho}=(\rho,\overline{\rho},1,\cdots,1)$, we see that $\langle\rho,x\rangle\langle\overline{\rho},x\rangle\langle 1,x\rangle^{n-1}=|\langle\rho,x\rangle|^{2}\langle 1,x\rangle^{n-1}=|\langle\rho,y\rangle|^{2}\langle 1,y\rangle^{n-1}$ and therefore $|\langle\rho,x\rangle|=|\langle\rho,y\rangle|$. Note that here we relied on the fact that $x$ and $y$ are real. This implies that the following map $\eta:\ G^{\vee}\rightarrow\mathbb{C}$ takes values in $\textnormal{U}(\mathbb{C})$:
$\eta(\rho)=\frac{\langle\rho,x\rangle}{\langle\rho,y\rangle}.$ (46)
Now, $\eta(1)=1$ since $\langle 1,y\rangle=\langle 1,y\rangle$. By setting $\underline{\rho}=(\rho,\mu,\overline{\rho\mu},1,\cdots,1)$ we see that $\eta(\rho)\eta(\mu)=\eta(\overline{\rho\mu})$ for all $\rho,\mu\in G^{\vee}$ and therefore $\eta\in(G^{\vee})^{\vee}$. Since the Fourier transform sends $G^{\vee}\subseteq\langle G\rangle$ to $(G^{\vee})^{\vee}\subseteq\langle G^{\vee}\rangle$, there exists $g\in G$ such that $\eta(\rho)=\overline{\rho(g)}$. This means that $\langle\rho,x\rangle=\overline{\rho(g)}\langle\rho,y\rangle$, which implies $x=g\cdot y$ by the equivariance properties of the Fourier transform. ∎
Spectral invariants can be defined in the non-commutative case, but arise subtleties. First, we replace the group structure of $G^{\vee}$ with the tensor product $\otimes$ of unitary representations. However, $\textnormal{Irr}(G)$ is not closed with respect to $\otimes$. This is circumvented by considering Clebsch-Gordan coefficients, i.e. irreducible unitary sub-representations of tensor products. This leads to the following definition of operator-valued spectra.
Definition 11.
Fix $n>0$ and $\underline{\rho}=(\rho_{V_{1}},\cdots,\rho_{V_{n}})\in\textnormal{Irr}(G)^{n}$. The spectrum of order $n$ associated to $\underline{\rho}$ is defined for $x\in\langle G\rangle$ as:
$\beta_{\underline{\rho}}(x)=\hat{x}_{\rho_{V_{1}}}\otimes\cdots\otimes\hat{x}_{\rho_{V_{n}}}\left(\hat{x}_{\rho_{T_{1}}}^{\dagger}\oplus\cdots\oplus\hat{x}_{\rho_{T_{k}}}^{\dagger}\right)\ \in\textnormal{End}(V_{1}\otimes\cdots\otimes V_{n})$ (47)
@@
The completeness of spectra of order $n\geq 2$ (Proposition B.2) extends to the non-commutative case [26].

### v3

@@ Appendix B Spectral Invariants
In this section, we overview the theory of invariants over $\langle G\rangle$, i.e. (polynomial) maps $\langle G\rangle\rightarrow\mathbb{C}$ that are invariant with respect to the action by $G$. To this end, we recall the following notion for a commutative group $G$.
Definition B.1.
@@
$\beta_{\underline{\rho}}(x)=\hat{x}_{\rho_{1}}\cdots\hat{x}_{\rho_{n}}\overline{\hat{x}}_{\rho_{1}\cdots\rho_{n}}.$ (36)
The spectra of order $n$ are invariant polynomials of degree $n+1$ containing one conjugate variable. The presence of the latter is necessary for invariance. For $n=1,2$ they are alternatively referred to as power spectra and bispectra respectively. Note that the power spectra reduce simply to $\beta_{\rho}(x)=|\hat{x}_{\rho}|^{2}$, $\rho\in G^{\vee}$, and constitute a standard tool in signal processing. Bispectra, together with higher-order spectra, were first introduced by Kakarala (2012). It is immediate to see that spectra generate all the polynomial invariants of $\langle G\rangle$ – see also (Sturmfels, 2008, Theorem 2.1.4).
@@
Proof B.3.
This follows from interpreting the invariance condition via the Fourier transform. Namely, given $\underline{\rho}\in(G^{\vee})^{n+1}$ consider the monomial over $\langle G^{\vee}\rangle$ defined by $\hat{x}_{\rho_{1}}\cdots\hat{x}_{\rho_{n}}\overline{\hat{x}}_{\rho_{n+1}}$. Since $g\cdot\hat{x}=(\overline{\rho(g)}\ \hat{x}_{\lambda})_{\lambda\in G^{\vee}}$, the monomial is invariant if and only if $\rho_{1}(g)\cdots\rho_{n}(g)\overline{\rho}_{n+1}(g)=1$ for all $g\in G$, i.e. $\rho_{n+1}=\rho_{1}\cdots\rho_{n}$. Since monomials linearly generate polynomials, the claim follows.
A remarkable aspect of spectra of even order is the fact that they jointly determine real (generic) elements of $\langle G\rangle$ up to the action by $G$ – a property known as completeness. This was first shown by Smach et al. (2008). For convenience, we report below a simple proof for finite commutative groups.
@@
Proof B.5.
By setting $\underline{\rho}=(1,\ldots,1)$ we see that $\beta_{\underline{\rho}}(x)=\hat{x}_{1}^{n+1}=\beta_{\underline{\rho}}(y)=\hat{y}_{1}^{n+1}\in\mathbb{R}\setminus\{0\}$ and therefore $\hat{x}_{1}=\hat{y}_{1}$ since $n$ is even. For $\rho\in G^{\vee}$, by setting $\underline{\rho}=(\rho,\overline{\rho},1,\ldots,1)$, we see that $\hat{x}_{\rho}\ \hat{x}_{\overline{\rho}}\ \hat{x}_{1}^{n-1}=|\hat{x}_{\rho}|^{2}\ \hat{x}_{1}^{n-1}=|\hat{y}_{\rho}|^{2}\ \hat{y}_{1}^{n-1}$ and therefore $|\hat{x}_{\rho}|=|\hat{y}_{\rho}|$. Note that here we relied on the fact that $x$ and $y$ are real. This implies that the following map $\eta:\ G^{\vee}\rightarrow\mathbb{C}$ takes values in $\textnormal{U}(\mathbb{C})$:
$\eta(\rho)=\frac{\hat{x}_{\rho}}{\hat{y}_{\rho}}.$ (37)
Now, $\eta(1)=1$ since $\hat{x}_{1}=\hat{y}_{1}$. By setting $\underline{\rho}=(\rho,\mu,\overline{\rho\mu},1,\ldots,1)$ we see that $\eta(\rho)\eta(\mu)=\eta(\overline{\rho\mu})$ for all $\rho,\mu\in G^{\vee}$ and therefore $\eta\in(G^{\vee})^{\vee}$. Since the Fourier transform sends $G^{\vee}\subseteq\langle G\rangle$ to $(G^{\vee})^{\vee}\subseteq\langle G^{\vee}\rangle$, there exists $g\in G$ such that $\eta(\rho)=\overline{\rho(g)}$. This means that $\hat{x}_{\rho}=\overline{\rho(g)}\ \hat{y}_{\rho}$, which implies $x=g\cdot y$ by the equivariance properties of the Fourier transform.
Spectral invariants can be defined in the non-commutative case, but give rise to subtleties. First, we replace the group structure of $G^{\vee}$ with the tensor product $\otimes$ of unitary representations. However, $\textnormal{Irr}(G)$ is not closed with respect to $\otimes$. This is circumvented by considering Clebsch-Gordan coefficients, i.e. irreducible unitary sub-representations of tensor products. This leads to the following definition of operator-valued spectra.
Definition B.6.
Fix $n>0$ and $\underline{\rho}=(\rho_{V_{1}},\ldots,\rho_{V_{n}})\in\textnormal{Irr}(G)^{n}$. The spectrum of order $n$ associated to $\underline{\rho}$ is defined for $x\in\langle G\rangle$ as:
$\beta_{\underline{\rho}}(x)=\hat{x}_{\rho_{V_{1}}}\otimes\cdots\otimes\hat{x}_{\rho_{V_{n}}}\left(\hat{x}_{\rho_{T_{1}}}^{\dagger}\oplus\cdots\oplus\hat{x}_{\rho_{T_{k}}}^{\dagger}\right)\ \in\textnormal{End}(V_{1}\otimes\cdots\otimes V_{n}),$ (38)
@@
The completeness of spectra of order $n\geq 2$ (Proposition B.4) extends to the non-commutative case (Kakarala, 2009). The proof we provide of Proposition B.4 translates almost identically to non-commutative groups, assuming $x,y\in\mathbb{R}^{G}$ have non-singular Fourier coefficients, i.e., $\hat{x}_{\rho_{V}},\hat{y}_{\rho_{V}}\in\textnormal{GL}(V)\subseteq\textnormal{End}(V)$. The only substantial difference in the proof is that, in order to conclude that $\eta(\rho)=\rho(g)^{\dagger}$ for some $g\in G$, it is necessary to invoke Tannaka-Krein duality.